rvc (rvc):

the extremity of the hour hand of a clock is 1/20th as fast as the minute hand .What is the length of the hour hand if the minute hand is 10cm long?

rvc (rvc):

@texaschic101 @praxer @radar

OpenStudy (anonymous):

Hour hand goes 360o in 12 hours or 30o/hr Minute hand goes 360o/hr = 12x hour hand To get it to be 20x speed of hour hand need shorter hour hand = (12/20)(10) = 6 cm long